OpenStudy (anonymous):

2.The total number of degrees in a regular hexagon is 720°. If a corresponding angle in a congruent regular hexagon is (11x – 12)°, find the value of x.

OpenStudy (anonymous):

Each angle in a regular hexagon is: 720/6 = 120 Set 120 = 11x - 12 and solve for "x"

OpenStudy (anonymous):

120=11x-12 +12 +12 132=11x 132/11=11x/11 12=x

OpenStudy (anonymous):

You got it! Good job!

OpenStudy (anonymous):

All good now, @Kaylei104 ?

OpenStudy (anonymous):

Yea. thank you

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Kaylei104

OpenStudy (anonymous):

Thank you and you're welcome